OpenStudy (anonymous):

The fertilizer ammonium sulfate, (NH4)2SO4, is prepared by the reaction between ammonia (NH3) and sulfuric acid: 2NH3 + H2SO4 --> (NH4)2SO4 How many kilograms of NH3 are needed to produce 1.40 x 10^5 kg of (NH4)2SO4?

OpenStudy (anonymous):

you will need half the amount of nh3 beacuse for every 2 mols of nh3 there is one mole of (NH4)2(SO4) so you just divide it by two.

OpenStudy (anonymous):

so you will ned 7*10^4

OpenStudy (anonymous):

I'm not following what you are saying.

OpenStudy (anonymous):

the equation states that for every 2 moles of nh3 you will get one mole of (nh4)2so4

OpenStudy (anonymous):

do you know how i got that?

OpenStudy (anonymous):

Not exactly

OpenStudy (anonymous):

I went and did what you did getting the 7 x 10^4 but it said it was wrong. That's more where I'm confused. I think I'm seeing where you are getting the numbers. It's from the chemical equation right?

OpenStudy (anonymous):

yes and sorry i did it in a hurry earlier see what you have to do is convert 1.4*10^5 kg of (nh4)2s04 to kg of nh3 1.4*10^5(NH4)2SO4

OpenStudy (anonymous):

\[\frac{ 1.4*10^5 kg NH4 }{ }*\frac{ 1000g NH4 }{ 1 kg NH4}*\frac{ 1 mole NH4 }{ 132g (NH4)2SO4 }*\frac{ 2 mole NH3 }{ 1 mol (NH4)2SO4 }*\frac{17 g NH3 }{ 1 mole NH3}*\frac{ 1 kg NH3}{ 1000 gNH3 }\]

OpenStudy (anonymous):

I got 36060.60606

OpenStudy (anonymous):

\[\frac{ 17 g }{ 1 mole nh3}*\frac{ 1 kg NH3 }{ 1000 g}\]

OpenStudy (anonymous):

that is the part that didnt fit so divide what you got by 1000 to get kg

OpenStudy (anonymous):

36.06

OpenStudy (anonymous):

Or would it be .00613

OpenStudy (anonymous):

I don't think I'm getting the right answer.

OpenStudy (anonymous):

it is the correct answer 3.61*10^4 put it in scientific notation

OpenStudy (anonymous):

I see now

OpenStudy (anonymous):

Thank you for the help!

OpenStudy (anonymous):

Would you be willing to help with another question?

OpenStudy (anonymous):

sure